OpenStudy (anonymous):

f(x) = x4 + 64x2 f(x) = x2(8x + i)(8x - i) f(x) = x2(x + 8i)2 f(x) = x2(x + 8i)(x - 8i) f(x) = x2(8x + i)2

OpenStudy (anonymous):

f(x) = x2(x + 8i)(x - 8i)

OpenStudy (anonymous):

thats what I got

OpenStudy (mathstudent55):

f(x) = x4 + 64x2 = x^2(x^2 + 64) = x^2(x^2 - (-64) ) = x^2(x + sqrt(-64) )(x - sqrt(-64) ) = x^2(x + 8i)(x - 8i)

OpenStudy (radar):

I believe you are on the right track. Congrats.

OpenStudy (anonymous):

Thank you guys so much!

OpenStudy (mathstudent55):

This is more elegant: f(x) = x4 + 64x2 = x^2(x^2 + 64) = x^2(x^2 - (-64) ) = x^2(x^2 - (8i)^2) Now you see it's a difference of two squares. = x^2(x + 8i)(x - 8i)

OpenStudy (anonymous):

yep just got a 100 on my math quiz, thanks guys

OpenStudy (mathstudent55):

wlcm